solve by factoring: 2x^2+3x=2 ...please. thanks!
Answer by stanbon(75887)

2x^2+3x=2
----
2x^2 + 3x -2 = 0
2x^2 +4x-x -2 = 0
2x(x+2) -(x+2) = 0
(x+2)(2x-1) = 0
-------------
x = -2 or x = 1/2
=====================
Cheers,
Stan H.